• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 00:19:54 on April28,2003

    Entry number 99029

    Followups:

    keyword=Magnet Settings

    Quads are back on and have been cycled. Septum set to Iset=152.8, Iout=168.75.
    Septum settings determined from quad-off studies. SNAKE indicates that we expect to run out of acceptance (with quads off) at 25% above the correct septum setting. We observed the elastic peak essentially fading away at 75% above what we have heretofore called the "nominal" setting. Inspection of th_rot vs x_rot spectra for different septum settings suggests that Iset-152.8 is about right. Thin sieve has been installed and we shall see.


    FIGURE 1

    FIGURE 2